# The Essence of Contextual Understanding in Theory of Mind: A Study on Question Answering with Story Characters

### v1

@@ 6.1 Setup
We conduct experiments on our generative and multichoice QA tasks in English and Chinese. Models are instructed to respond with vanilla prompts as in Appendix I. To investigate the capability of the model to exploit contextual information for ToM comprehension, we vary the context lengths of story plots exposed to models.
@@ 6.3 Human Performance
w/. history w/o. history GPT-4o o1
c=0 c=2k c=0 c=2k c=0 c=2k c=0 c=2k
all 59.3 62.0 41.3 48.7 47.3 42.0 51.3 43.3
indir. 54.8 60.6 33.7 40.4 45.1 37.5 46.2 39.4
Table 3: Accuracy of multichoice QA (Chinese) by different groups of participants who are familiar or unfamiliar with selected books. We compare with GPT-4o and o1 on the same subset. indir. refers to indirect questions.
@@ 6.4 Analysis and Discussions
Can CoT (o1 models) help?
We further explore whether CoT prompting could bring improvement of ToM comprehension. For this study, we use the most advanced OpenAI o1 model. Because of the high complexity and cost of o1-model, we compare it on the human study subset of 150 instances. As shown in Table 3, o1-model improves over GPT-4o, showing the benefit from stronger reasoning ability. However, it still lags 10% behind humans.
@@
This failure can be attributed to two potential reasons. First, despite the ability to process long inputs, LLMs struggle with capturing dependencies between paragraphs and enhancing the understanding of individual paragraphs based on these dependencies Xu et al. (2024). Second, the LLMs may solve our task largely with their memory of the books and reviews, exhibiting the Clever Hans phenomenon Shapira et al. (2024) – as the evaluated books and their reviews have been exposed to the LLMs during pre-training Jiang and Ferraro (2024). It is likely that the LLMs can recall the associated plot details and knowledge with both short and long windows. We leave this study of these reasons in future work.
@@
As expected, both humans unfamiliar with the books and the LLMs drop dramatically on the subset of indirect questions. Furthermore, it is noteworthy that the o1 model shows limited improvement on this subset, despite its superior performance on the full set. It suggests that o1 does well mainly via exploring the superficial textual clues between answers and plot texts, while its ToM reasoning capability is still limited.
@@ 7 Conclusion
We introduce CharToM-QA benchmark that aims to evaluate machines’ ToM capability in realistic scenarios with complex social relationships and interactions. The benchmark focuses on the aspect of contextual understanding about the complex background stories, which is largely igonored in existing benchmarks. The tasks in the benchmark take the forms of generative and multichoice QA. Particularly, for generative QA, we design an evaluation protocol that inspects the bonus points and penalty existing in model responses. Experimental results reveal that current advanced LLMs, including o1, still lag behind humans and struggle at exploiting context to enhance the understanding about the ToM of characters.
@@
The construction of CharToM-QA benchmark involves the usage of public books and notes written by reading app users. During the process of using these data, we strictly adhere to the guidelines in Internet Research Ethics. All the books are available in the Gutenberg project and their Chinese-translated version are properly used with valid license. For the notes, we collect and use them with necessary procedures to protect the privacy of online reading app users who wrote the notes. For annotators participating in the benchmark construction, they are adequately paid according to their working hours.
@@ Appendix A Book List
We keep the books with top 20 user notes for the construction of CharToM-QA benchmark. Table 4 lists the chosen books and the number of questions belonging to each book in the descending order.
@@
Appendix B Definition of ToM Dimensions
@@
Appendix C Statistics of Questions and Bonus Points.
@@
Appendix D Prompt for Book Note Filtering
@@
Appendix E Prompt for ToM Paraphrasing
@@
Appendix F Prompt for Question Generation & Verification
@@
Assuming you are an expert in psychology and literary. You are reading the book \${book_name}. There are a piece of description about the \${dimension} of \${target_character} and a question related to the book story. Please judge whether [Candidate Question] is an appropriate question whose direct answer is [ToM description]. [Candidate Question]: \${candidate_question} [ToM description]: \${ToM_description} An appropriate question should satisfy the following requirements: (1) [Candidate Question] should not be too broad. (2) [ToM description] should be a fluent and logically correct answer to [Candidate Question]. Please answer in the following format: “Appropriate / Inappropriate” “Reason”
@@
Appendix G Prompt for Extracting Bonus Points
@@
Appendix H Construction of Difficult Distraction Choices for Multichoice QA
@@
Appendix I Prompt for Responding to Questions
Table 12 and Table 13 give the prompts for responding to the questions in generative QA and multichoice QA during experiments, respectively.
@@ Appendix I Prompt for Responding to Questions
Assuming you are an expert in psychology and literary. Please construct appropriate multichoice questions for the given [Story Plot] from the the book \${book_name}. [Story Plot]: \${story_plot} [Question]: \${question} [Reference Answer]: \${reference_answer} [Misleadings]: \${misleadings} Note that (1) Ensure that the reference answer is the only correct answer to the question. (2) The three distraction candidate choices should be seemingly plausible but actually unreasonable. (3) You can refer to the reference answer and misleadings to construct distraction choices. (4) Do not repeat the same content in different distraction choices. Please provide distraction candidate choices that meet the above requirements and conform to the following format: [Distraction Choice 1]: (Choice Content) [Distraction Choice 2]: (Choice Content) [Distraction Choice 3]: (Choice Content)
@@
Assuming you are an expert in psychology and literary. Based on your profound understanding of the story in the book \${book_name}, Please answer the [Question]. The following items give the [Story Plot] and the [Question]. [Story Plot]: \${story_plot} [Question]: \${question} Note that you should respond to the question using only one concise sentence, with around \${length_of_answer} tokens. Your response is:
@@
Assuming you are an expert in psychology and literary. Based on your profound understanding of the story in the book \${book_name}, Please answer the [Question]. The following items give the [Story Plot], the [Question] and the four candidates from [Candidate Choices].. [Story Plot]: \${story_plot} [Question]: \${question} [Candidate Choices]: (1). \${cand_choice_1} (2). \${cand_choice_2} (3). \${cand_choice_3} (4). \${cand_choice_4} Note that 1. You should only choose one candidate from (1),(2),(3),(4). 2. Only output the index of your chosen candidate. 3. Do not include any other unnecessary content or symbol. Your choice is:
@@
Appendix J Prompt for Generative QA Evaluation
@@
Appendix K Illustrative Cases
@@
…… Bring us some liquor, Turk!’“Here it is, sir,” said Christophe, holding out the bottle.Vautrin filled Eugene’s glass and Goriot’s likewise, then he deliberately poured out a few drops into his own glass, and sipped it while his two neighbors drank their wine.All at once he made a grimace.“Corked” he cried. ……
@@
Bonus Point:
 (1) Foutriquet deliberately tries to please Eugène de Rastignac and Monsieur Goriot.
 (2) To ensure that his “business” does not go sour.
@@
 (3) Vautrin deliberately flatters Eugène and Father Goriot to ensure that his b̈usinessd̈oes not fall through.

### v2

@@ 6.1 Setup
We conduct experiments on our generative and multichoice QA tasks in English and Chinese. Models are instructed to respond with vanilla prompts as in Appendix K. To investigate the capability of the model to exploit contextual information for ToM comprehension, we vary the context lengths of story plots exposed to models.
@@ 6.3 Human Performance
w/. history w/o. history GPT-4o o1 DS-R1
c=0 c=2k c=0 c=2k c=0 c=2k c=0 c=2k c=0 c=2k
all 59.3 62.0 41.3 48.7 47.3 42.0 51.3 43.3 54.0 48.7
indir. 54.8 60.6 33.7 40.4 45.1 37.5 46.2 39.4 50.4 40.2
Table 3: Accuracy of multichoice QA (Chinese) by different participants familiar or unfamiliar with selected books. We compare with GPT-4o, o1 and DS-R1 on the same subset. indir. refers to indirect questions.
@@ 6.4 Analysis and Discussions
Can Chain-of-Thought (CoT) help?
We further explore whether CoT could bring improvement of ToM comprehension. For this study, we use the recently released strong OpenAI o1 and DeepSeek-R1 (DS-R1) Guo et al. (2025) models on the human study subset of 150 instances. As shown in Table 3, both o1 and DS-R1 outperform GPT-4o, showing the benefit from stronger reasoning ability. However, they still notably lag behind humans. As the CoT of DS-R1 is visible, we closely examine its generated thoughts on randomly selected 100 samples. We find that 76% of the thoughts initially recalled either the character’s personal background or prior plot events that are relevant to the question. Among the remaining 24% of cases, three-fourths of questions are about emotions, which is as expected because emotions are more likely to be temporary mental states that rely less on global contexts. These findings further highlights the importance of comprehensive contextual understanding for ToM reasoning, a promising direction for enhancing future LLM.
@@
This failure can be attributed to two potential reasons. First, despite the ability to process long inputs, LLMs struggle with capturing dependencies between paragraphs and enhancing the understanding of individual paragraphs based on these dependencies Xu et al. (2024b). Second, the LLMs may solve our task largely with their memory of the books and reviews, exhibiting the Clever Hans phenomenon Shapira et al. (2024) – as the evaluated books and their reviews have been exposed to the LLMs during pre-training. It is likely that the LLMs can recall the associated plot details and knowledge with both short and long windows. We leave this study of these reasons in future work.
@@
As expected, both humans unfamiliar with the books and the LLMs drop dramatically on the subset of indirect questions. Furthermore, it is noteworthy that both o1 and DS-R1 models also perform worse on the indirect questions than on the full set. As indirect questions typically require broader global context of the whole stories rather than only superficial textual clues in the given plot windows, it suggests that the ToM capabilities of current LLMs are still limited especially in the aspect of comprehensive contextual understanding.
@@ 7 Conclusion
We introduce CharToM-QA benchmark that aims to evaluate machines’ ToM capability in scenarios with complex social relationships and interactions. CharToM-QA focuses on the aspect of contextual understanding in comprehending ToM of characters in novels, which is largely ignored in existing benchmarks. The tasks take the forms of generative and multichoice QA. For generative QA, we design an evaluation protocol inspecting bonus points and penalty in model responses. Experimental results reveal that current advanced LLMs, including o1 and DS-R1, still lag behind humans and struggle at effectively exploiting context to comprehend ToM of characters. Overall, our CharToM-QA benchmark highlights the importance of comprehensive contextual understanding for advancing ToM reasoning in LLMs.
@@
The construction of CharToM-QA benchmark involves the usage of public books and notes written by reading app users. During the process of using these data, we strictly adhere to the guidelines in Internet Research Ethics. All the books are available in the Gutenberg project and their Chinese-translated version are properly used with valid license. For the notes, we collect and use them with necessary procedures to protect the privacy of online reading app users who wrote the notes. For annotators participating in the benchmark construction, they are adequately paid according to their working hours. In our human study, we do not collect any personally identifiable information of the participants. The participants are only asked to respond to multichoice questions and not required to provide any new written input.
@@ Appendix A Book List
We keep the books with top 20 user notes for the construction of CharToM-QA benchmark. These books are publicly accessible and span different historical periods, genres, and topics. Table 4 lists the chosen books and the number of questions belonging to each book in the descending order. It is also notable that our dataset has included a number of unique target characters (150 in total) from these books. These characters are significantly diverse, crafted by different authors with varying literary styles. Particularly, these characters possess distinct personalities, encompassing various types and fundamentally different personal backgrounds. These elements could introduce diversity to our dataset, effectively mitigating potential biases related to literary style and topic.
@@
Appendix B ToM Dimensions Studied
@@
Benchmarks w/. Template-based Problem Construction
Task Form Story Creation Problem Construction Character Background #Eval_Q
Template AI Human Template AI Human
ToMi Nematzadeh et al. (2018) MQA N/A 6,000
Social IQA Sap et al. (2019) MQA N/A 2,224
Hi-ToM Wu et al. (2023) MQA N/A 1,000
FanToMKim et al. (2023) MQA,GQA N/A 2,118
BigToM Gandhi et al. (2023) MQA N/A 5,000
NegotiationToM Chan et al. (2024) MQA pre-defined character’s preference items 13,800
OpenToM Xu et al. (2024a) MQA pre-defined personlity traits and preferences 16,008
COKE Wu et al. (2024) NLG N/A -
Benchmarks w/o. Template-based Problem Construction
Task Form Story Creation Problem Construction Character Background #Eval_Q
Template AI Human Template AI Human
FauxPas-EAI Shapira et al. (2023) MQA,GQA N/A 80
ToMChallenges Ma et al. (2023a) MQA,GQA N/A 360
ToMBench Chen et al. (2024) MQA N/A 2,860
CharToM-QA (Ours) MQA,GQA all of the character’s previous experience depicted in story 1,035
Table 5: Comparison among CharToM-QA and other ToM benchmarks in terms of task form, story generation, problem construction, character background and the number of evaluation questions (#$\#Eval\_Q$). In the column of “Task Form”, MQA, GQA and NLG refer to multichoice QA, generative QA and natural language generation tasks, respectively. “✓” denotes that a benchmark possesses corresponding properties. “N/A” means the story merely depicts character’s current situation without extra character background. “-” in the #$\#Eval\_Q$ column means the task form does not support the counting of questions equivalent to standard QA settings. *The COKE and Social IQA datasets are essentially about thoughts on input events/situations so the concept of stories does not apply.
Appendix C Comparison of Benchmark Characteristics
In Table 5, we compare the characteristics of our CharToM-QA and other ToM benchmarks. First, our CharToM-QA provides both multichoice QA and generative QA tasks. Second, as the classic novel stories are directly used, CharToM-QA needs no additional story creation procedure, unlike most of other benchmarks. Meanwhile, it naturally possesses high-quality diverse social scenarios with complex social relationships and interactions due to the intrinsic features of human-written novels. Third, for the problem construction, most other benchmarks involve templates for generating questions or answers. Although this kind of template-based approach can produce a relatively large number of problems, it could bring shortcuts and spurious correlations into benchmarks. In contrast, we adopt an AI-assisted human annotation strategy elaborated in Section 4 without any pre-defined template. This simultaneously ensures the efficiency and quality of our annotation process.
Last and most importantly, in the aspect of character background, CharToM-QA is the only benchmark that incorporates character’s previous experience depicted in story. Most of other ToM benchmarks merely provide depictions of character’s current situation or some pre-defined background information such as conversation topics, preference items. As demonstrated by the human study described in Section 6.3, such personal background is necessary for comprehensive contextual understanding about characters to correctly answer ToM-related questions in CharToM-QA. Specifically, participants with similar educational levels who have read the book (and thus possess knowledge of long personal backgrounds) perform significantly better than those who have not read when given the same contexts. It shows that our CharToM-QA highlights the importance of comprehensive contextual understanding for ToM reasoning, an aspect overlooked by previous benchmarks.
@@ Appendix C Comparison of Benchmark Characteristics
Appendix D Statistics of Questions & Bonus Points
@@
Appendix E Correlations of Evaluation Metrics with Humans on Generative QA
For Generative QA, besides the designed LLM-based evaluation protocol, we also use conventional token-based (RougeL) and embedding-based (BertScore) metrics for assessing the quality of model responses. In Table 7, we report the experimental results in terms of BPC, RougeL and BertScore when the context length is 2k. The reported values are the weighted averages of corresponding metrics over the four ToM dimensions.
BPC(%) RougeL(%) BertScore(%)
GPT-4o 39.8 21.5 56.6
GPT-3.5-Turbo-1106 34.0 19.6 51.6
Llama-3.1-8B-Instruct 33.5 16.2 50.8
Qwen2-7B-Instruct 34.9 15.9 50.5
Mistral-7B-Instruct-v0.3 29.9 19.0 49.7
InternLM2-7B-Chat 27.6 17.7 49.7
Human Correlation* 0.92 0.72 0.81
Table 7: Generative QA performances of LLMs in terms of BPC, RougeL and BertScore with the context length being 2k. “*”: The human correlation of these metrics are calculated on the subset used in our human study.
From the table, our LLM-based metric gives similar relative rankings among different LLMs as conventional metrics. Moreover, on the subset used in our human study, we let human experts assess model responses by giving scores ranging from 0 (nonsense) to 1.0 (perfect). The Pearson Correlations of human judgments with BPC, RougeL and BertScore are 0.92, 0.72 and 0.81, respectively. These results further justify the reliability of our designed evaluation protocol.
Appendix F Prompt for Book Note Filtering
@@
Appendix G Prompt for ToM Paraphrasing
@@
Appendix H Prompt for Question Generation & Verification
@@
Assuming you are an expert in psychology and literary. You are reading the book \${book_name}. There are a piece of description about the \${dimension} of \${target_character} and a question related to the book story. Please judge whether [Candidate Question] is an appropriate question whose direct answer is [ToM description].
[Candidate Question]: \${candidate_question}
[ToM description]: \${ToM_description}
An appropriate question should satisfy the following requirements: (1) [Candidate Question] should not be too broad. (2) [ToM description] should be a fluent and logically correct answer to [Candidate Question].
Please answer in the following format:
“Appropriate / Inappropriate”
“Reason”
@@
Appendix I Prompt for Extracting Bonus Points
@@
Appendix J Construction of Difficult Distraction Choices for Multichoice QA
@@
Appendix K Prompt for Responding to Questions
Table 14 and Table 15 give the prompts for responding to the questions in generative and multichoice QA during experiments, respectively.
Assuming you are an expert in psychology and literary. Based on your profound understanding of the story in the book \${book_name}, extract the critical bonus points from the [Answer] for the [Question]. The bonus points are used to comprehensively assess the quality of responses from other respondents. Below are the [Story Plot], [Question] about the character \${target_character} and the [Reference Answer].
[Story Plot]: \${story_plot}
[Question]: \${question}
[Reference Answer]: \${reference_answer}
Note that (1) Bonus points must be derived from the reference answer without hallucination. Bonus points are the indispensable points in the reference answer, but should not include any unnecessary explanatory content. (2) Different bonus points should orient to different aspects of the reference answer. (3) The semantic granularity of the bonus points should not be too fine. Each bonus point must be a sentence that constitutes a complete semantic unit. Semantically coherent content in the reference answer should not be broken down into multiple bonus points.
@@
Assuming you are an expert in psychology and literary. Please construct appropriate multichoice questions for the given [Story Plot] from the the book \${book_name}.
[Story Plot]: \${story_plot}
[Question]: \${question}
[Reference Answer]: \${reference_answer}
[Misleadings]: \${misleadings}
Note that (1) Ensure that the reference answer is the only correct answer to the question. (2) The three distraction candidate choices should be seemingly plausible but actually unreasonable. (3) You can refer to the reference answer and misleadings to construct distraction choices. (4) Do not repeat the same content in different distraction choices.
Please provide distraction candidate choices that meet the above requirements and conform to the following format:
[Distraction Choice 1]: (Choice Content)
[Distraction Choice 2]: (Choice Content)
[Distraction Choice 3]: (Choice Content)
@@
Assuming you are an expert in psychology and literary. Based on your profound understanding of the story in the book \${book_name}, Please answer the [Question]. The following items give the [Story Plot] and the [Question].
[Story Plot]: \${story_plot}
[Question]: \${question}
Note that you should respond to the question using only one concise sentence, with around \${length_of_answer} tokens.
Your response is:
@@
Assuming you are an expert in psychology and literary. Based on your profound understanding of the story in the book \${book_name}, Please answer the [Question]. The following items give the [Story Plot], the [Question] and the four candidates from [Candidate Choices]..
[Story Plot]: \${story_plot}
[Question]: \${question}
[Candidate Choices]:
(1). \${cand_choice_1}
(2). \${cand_choice_2}
(3). \${cand_choice_3}
(4). \${cand_choice_4}
Note that
1. You should only choose one candidate from (1),(2),(3),(4).
2. Only output the index of your chosen candidate.
3. Do not include any other unnecessary content or symbol.
Your choice is:
@@
Appendix L Prompt for Generative QA Evaluation
@@
Appendix M Illustrative Cases
@@
…… Bring us some liquor, Turk!” “Here it is, sir,” said Christophe, holding out the bottle. Vautrin filled Eugene’s glass and Goriot’s likewise, then he deliberately poured out a few drops into his own glass, and sipped it while his two neighbors drank their wine. All at once he made a grimace. “Corked” he cried. ……
@@
Question: Why did Vautrin choose to fill Eugène and Father Goriot’s glasses first?
@@
Bonus Points:
 (1) Vautrin deliberately tries to please Eugène and Father Goriot.
 (2) To ensure that his “business” does not fall through.
@@
 (3) Vautrin deliberately flatters Eugène and Father Goriot to ensure that his “business” does not fall through.